OpenStudy (s):

Calculate the volume of ).1 M NaOH solution you need to add to 50.0 ml of 0.1 M CH3COOH solution to form a buffer ph 4?

OpenStudy (abb0t):

I think you're missing some information here, you need the pKa of acetic acid. But anyway, you can use the henderson-hasselbach equation here to find the concentration and finally, get the volume. \(\sf \color{red}{pH = pK_a + log[\frac{A^-}{HA}]}\) \(\sf \color{blue}{pH-pKa = log(\frac{[A^-]}{[HA]})}\) \(\sf \color{purple}{10^{(pH-pK_a)}=\frac{[A^-]}{[HA]}}\) find your expression for [HA] = 0.005-x the 0.005 moles comes from: 0.1M\(\times\)(0.05 L) remember, M = \(\sf \frac{mol}{L}\)

OpenStudy (abb0t):

If you have K\(_a\), you can get pKa bu taking the -log(K\(_a\))